Suppose that the battery failure time. measured in hours, has a probability density
function given by:

f(x) =2/(x+1)), x>=0

The probability that the battery lasts longer than five hours is:
1
Expert's answer
2020-11-02T20:19:36-0500

Suppose that the battery failure time. measured in hours, has a probability density

function given by:


"f(x) = \begin{cases}
 \dfrac{2}{(x+1)^3}, x\geq 0 \\
 \ \ \ \ \ \ \ 0 \ \ \ \ \ \ \ \ x<0 
\end{cases}"


"P(X>5)=1-P(X≤5)=1-\displaystyle\int_{0}^5\dfrac{2}{(x+1)^3}dx="

"=1-\big[-\dfrac{1}{(x+1)^2}\big]\begin{matrix}
 5 \\
 0
\end{matrix}=1+\dfrac{1}{36}-1=\dfrac{1}{36}"